[Brahms-l] Shift report 20050218 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Feb 18 2005 - 16:01:35 EST
Shifter leader was jth
Shifters were 


1. Beam from 9:10 until the fire alarm around 10:00. New beam 12:15-14:45.
2. MRS@40deg@B2255, FS@4deg@A3450
run     tr2      tr3
13235   151k      77k
13241   257k     130k
Non-TPC pedestal run: 13236, 13242
TPC pedestal run: 13240, 13243
3. Found a missing channel, see E-log

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 18 16:03:09 2005

This archive was generated by hypermail 2.1.8 : Fri Feb 18 2005 - 16:03:24 EST